# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ 4.2 Open-sourced LLMs for agentic applications
Domain Tools Studies
Knowledge Grounding, Web Search, and Structured Retrieval Bing Search API, Google Search, DuckDuckGo, Wikipedia API, PubMed, UMLS, ESI Handbook, FinBERT, MedRAG, ChemCrow (RoboRXN) [35, 66, 45, 92, 38, 102, 53, 39, 56]
Code Generation, API Use, and System-Level Integration Code Interpreters, AutoGen, Excel, Copilot, PowerBI, Jupyter AI, Chapyter, CoML, RDKit, Scikit-learn, RapidAPI, ChatEDA (OpenROAD), Speechly [35, 109, 34, 38, 69, 77, 66, 137, 52, 74, 114]
Interactive and Embodied Environments AI2THOR, ROS 2, Gazebo, Clearpath Husky, MineDojo (Minecraft), MarioAI, ALFWorld, SMAC, Overcooked-AI, DeCoAgent (Smart Contracts) [47, 103, 58, 122, 48, 13, 108, 25, 37]
This domain is further enriched by multimodal and architectural developments. Vision language assistants such as LLaVA-7B and the more recent LLaVA-v1.6-mistral-7B combine a CLIP-style image encoder with a chat-based LLM, allowing agent capabilities in embodied environments and GUI-based interfaces [133, 62]. CogVLM-17B expands this functionality with enhanced visual reasoning and understanding capacity [59]. In the realms of perception and generative synthesis tasks, agents commonly use OpenAI CLIP, Stable Diffusion XL for high resolution image generation, and VQ-GAN for producing latent image tokenization [58]. These components are often combined with the LLaMA-2 models in interactive applications.
These model ecosystems establish the foundational open source landscape on which LLM agent research is currently based.
@@ 5.1 Usage of tools across knowledge grounding, web search, and structured retrieval
In addition, these tools are often paired with structured external knowledge sources. In particular, agents operating in domain-specific settings. For example, biomedical agents are linked to PubMed and UMLS knowledge bases, one also found using the Emergency Severity Index (ESI) handbook [38, 38] for access to specialized medical information, FinBERT to classify sentiment of financial texts, while the Wikipedia Web API is a widely used source for general encyclopedic information [102, 45, 38, 92, 53].
@@ 5.2 Usage of tools across code generation, API use, and system-level integration
A widely used tool in agentic systems is the diverse types of Code Interpreter developed in a number of studies [35, 109]. This is particularly relevant for data science agents, which are evaluated for their integration with platforms and tools such as Excel, Copilot, Power BI, Jupyter AI, Chapyter and CoML [109]. Some studies also incorporate AutoGen, a framework designed to build agentic AI systems through multiagent interactions and improved LLM inference capabilities [34, 38, 69, 77]. These agents often produce code to interact with file systems [66], command-line utilities [66], and domain-specific libraries, such as RDKit for chemical informatics [137] and Scikit-learn for machine learning. Qin et al. found that using ChatGPT’s function calling mechanism, LLMs can interact with 16,464 RESTful APIs sourced from RapidAPI by utilizing a neural API retriever and structured API documentation [52].
In the Electronic Design Automation (EDA) sector, the ChatEDA agent demonstrates this capability through its Python wrapper-based interaction with the OpenROAD platform [74]. Cuadra et al. in their study used Speechly to support the speech-to-text task for the purpose of entering health data [114].
@@ 5.3 Usage of tools across interactive and embodied environments
To investigate complex and interactive behaviors, agents are often employed in simulated or real-world environments, which function as tools that offer detailed, state-dependent feedback in response to agent interactions. Zhu et al. [47] employed AI2THOR, a simulation platform that supports realistic 3D environments with physical interactions and dynamic visual states, for the purpose of assessing embodied agents within household environments.
In human-robot interaction, Frering et al. [103] utilized ROS 2 and the Gazebo simulator with a model like the Clearpath Husky. Interactive environments used in agentic research include virtual platforms like game worlds such as Minecraft (via MineDojo) [58, 122], the MarioAI platform [48], and text-based simulations such as ALFWorld [13]. In multiagent contexts, popular benchmarks include the StarCraft Multi-Agent Challenge (SMAC) and the collaborative game Overcooked-AI [108, 25]. An especially innovative example can be seen in the DeCoAgent framework, where LLM agents autonomously coordinate with the help of smart contracts, overcoming limitations of static, closed multi-agent environments [37].
@@ 6 Frameworks for building LLM agents
This section systematically reviews the frameworks developed for building LLM agents, highlighting their architectural designs, core capabilities, and operational domains. By categorizing existing solutions and analyzing their design principles, it provides a foundational lens for understanding how LLM agents are constructed and deployed in diverse scenarios.
@@ 6.1 Basic architecture of an LLM agent
Profiling. The profiling module defines an agent’s operational persona (e.g., developer, advisor or task-specific role), conditioning its behavior and role policy through static profiles defined by experts or dynamic generative mechanisms [21, 144]. Static profiles encode domain knowledge and role-specific behaviors, while dynamic profiles simulate human-like variability by generating diverse agent personas through prompt engineering or parameter sampling [106]. These profiles may include demographic data, personality traits, and social relationships, significantly influencing downstream decisions in memory retrieval, planning strategies, and action selection.
Memory. Memory is essential for enabling agents to remember past interactions, maintain coherence, and improve over time. To ensure temporal coherence, memory is typically divided into short-term (prompt-based) and long-term (externally stored) forms. Short-term memory supports in-the-moment reasoning by storing dialogue history and environmental signals, but is limited by LLM context windows [31]. Long-term memory captures reusable skills, patterns, or tools from past interactions, such as the skill libraries used in Voyager or the structured experience repositories employed in Reflexion [13].
Knowledge-augmented memory further extends the capability by retrieving relevant information from external sources. Memory formats vary from natural language, embeddings, databases (e.g., SQL), to structured lists, each chosen based on task requirements. Common operations include reading (prioritized by recency, relevance, and importance), writing (handling duplication and overflow), and reflection (summarizing past experiences into high-level insights).
Planning. The planning module decomposes complex tasks into intermediate goals and determines sequences of actions. This module breaks down complex goals into structured subtasks using strategies such as CoT [145], Tree-of-Thought (ToT) [146], or interaction-based feedback loops. Planning can occur with or without feedback. Feedback-free strategies often use stepwise prompting (e.g., CoT, ToT). Feedback-based iteration empowers agents to dynamically adapt plans using signals from the environment, human input, or reflections, often reinforced through techniques like Monte Carlo Tree Search or multi-agent collaboration.
@@
Together, these components enable LLMs to function as autonomous agents-reasoning, remembering, planning, and acting-in open-ended and evolving tasks. This modular architecture is foundational to both single-agent and multiagent systems discussed in the following sections.
@@ 6.1.1 Single-agent LLM system
LLM is the core of an agent, which allows task planning and decision making based on current inputs, memory, and feedback [148]. In addition, a objective defines the target goal and guides how the agent breaks down complex tasks to formulate a strategy. Memory is a dynamic storehouse of past interactions and relevant contextual information, enabling the agent to adapt its behavior thoughtfully [149, 150]. Action refers to the operational capabilities of the agent. It executes commands, interacts with tools or communicates outputs [14]; and Rethink facilitates reflective learning by evaluating previous actions and environmental feedback to inform future decisions [15].
@@
Figure 5: A general overview of a multi-agent LLM system. Here, three agents operate within a multimodal environment where they act, generate results, and exchange feedback. Each agent is equipped with internal modules (brain) including memory, reasoning, and planning, that guide its behavior. Through collaborative communication, agents perceive the environment, coordinate strategies, and ultimately take action.
@@ 6.1.2 Multi-agent LLM systems
Multi-agent LLM systems are designed for coordinated collaboration, where multiple agents communicate, adapt, and solve problems together (Figure 5). A description of how multiple agents observe, interact with, and adjust to different environments to work together toward shared objectives is given to highlight the foundation for effective multiagent collaboration.
@@
Communication structures. In this layer, three major paradigms are prevalent for agent communication, including cooperative (to work together toward shared goals), debate (to argue opposing views to reach consensus), and competitive (to pursue individual objectives). Communication can be centralized, decentralized, or shared memory based, where agents publish and subscribe to a shared message pool.
@@ 6.2 Common LLM agent frameworks
Several frameworks have been adopted to implement LLM agents, facilitating reasoning, decision making, memory management, and action execution in single-agent and multiagent contexts. Table 7 presents a categorized overview of these frameworks based on their application in single-agent or multi-agent systems.
Table 7: Common LLM agent frameworks used in single- and multi-agent LLMs
Agent Framework Type Ref.
ReAct Single [112, 91, 67, 102, 54, 113, 80, 110, 116, 68, 13, 100, 51]
Multi [90, 56]
Reflexion Single [91, 67, 102, 113, 115, 68, 13, 100, 51]
LangChain Single [106, 87, 116, 107, 119]
Multi [36, 63, 56]
AIDE Single [89]
AutoGen Multi [38, 36, 85, 69, 78, 77]
MetaGPT Single [109]
Multi [95, 36]
CAMEL Single [106]
Multi [36, 85, 22, 129]
Generative Agent Multi [63]
BDI Single [103]
Multi [129]
Among these, ReAct and Reflexion are the prominently utilized frameworks within single-agent LLM architectures, while AutoGen and CAMEL are widely employed in multi-agent environments. Furthermore, LangChain is recognized as a highly versatile framework extensively applied across single-agent and multi-agent paradigms.
@@ 6.3.1 Single-Agent LLM Systems’ Application Domain
Healthcare. LLM agents in healthcare have shown promise in tasks ranging from personalized consultation to domain-specific scientific analysis. Conversational agents provide personalized health advice and support elderly care through natural interfaces [80, 82]. Others optimize public health policies using reinforcement learning [107] or operate in real-world clinical settings [81]. More specialized agents help with cancer protein analysis and biomedical research tasks [79], underscoring the ability of single agent systems to operate in high-stakes and knowledge-intensive settings.
@@
General-purpose systems. A growing body of work seeks to develop general purpose agent frameworks that empower LLMs with reasoning, planning, and tool use capabilities [52, 112, 66, 53, 46]. These systems support multitask environments [91], enhance collaboration with humans [102], and enable long-horizon planning through world knowledge and instruction tuning [50, 100]. Several benchmarks have been proposed to assess performance in data science [109], safety [110], and reasoning tasks [49, 148], while architectures address self-improvement [98] and causal reasoning [126]. Reinforcement learning paradigms adapted to LLM further boost agent autonomy [68, 13, 130].
@@ 6.3.2 Multi-Agent LLM systems’ application domain
Table 8: Reasoning, planning, and memory techniques in LLM-based agents across single-agent, multi-agent, and widely-used categories
Reasoning techniques in LLM-based agents
Category Technique/Framework
Single-agent reasoning TOOLLLM depth-first search-based decision tree [52], EASYTOOL task decomposition [53], MATRIX simulation-based self-critique [98], ToolEmu risk-prompted reasoning [110], ToT [49], CogMir social-context prompts [105], Moral regret mechanisms [134], SHAP-based interpretability [64], Beneficial hallucination-driven hypothesis testing [119], ChatNT perplexity-informed confidence estimation [56]
Multi-agent reasoning Prompt-Structured Strategic Reasoning [108], AGENTVERSE multi-agent collaborative reasoning [35], Theory-of-Mind (ToM) inspired belief module [104]
Widely used techniques CoT [40, 104, 45, 47, 25, 110, 49, 105, 74, 63, 59, 117, 133, 62, 73, 121, 79], ReAct [13, 113, 102, 112, 91], Self-Reflection and Self-Critique [91, 19, 105, 60, 63, 117, 122, 13, 107, 133, 153, 111, 86, 83, 56], Self-Consistency methods [39, 54], Multi-Agent Collaborative Reasoning [19, 38, 69, 77], Ensemble and consensus strategies [39, 42, 100]
Planning techniques in LLM-based agents
Category Technique/Framework
Single-agent planning TIMEARENA heuristic planning [91], ReHAC multi-step planning [102], Task stack-based planning [47], Single-shot planning [124], World Knowledge Model guided planning [50], RAFA multi-step trajectory generation [135]
Multi-agent planning ProAgent modularized planning [25], Plan-first execution paradigm [87], DeCoAgent few-shot prompt-based task decomposition [37], Causal strategy learning systems [126], Star-Agents evolutionary strategy [44], UrbanKGent fixed pipeline [86], SMART-LLM coalition formation [61], Human-in-the-loop workflow [83]
Widely used techniques Multi-step planning [102, 124, 61, 135], Task decomposition planning [120, 37, 53, 91], ReAct-based planning [112, 91, 132], Reflexion planning [91, 132, 102], Re-planning and dynamic adaptation [118, 122], Plan generation and evaluation [97, 116]
Memory techniques in LLM-based agents
Category Technique/Framework
Single-agent memory RE-Bench [89], RCAgent’s OBSK [54], Memory buffers [114], VLN-Imagine [154], ToolEmu prompt context tracking [110], MLAgentBench compressed Research Logs [116], Formal Reasoning Graphs [49], STEVE-EYE episodic re-prompting [58], Moral alignment prompts [134], AGENTBENCH extended histories [148], CogAgent external screenshots and latent weights [59], Rolling logs for robots [103], Prior tool outputs in context [117], Negotiated ontologies [64], Concatenated dialogue contexts [122]
Multi-agent memory ProAgent recent-K and relevant-K retrieval [25], ColaCare retrieval-augmented generation [75], Satellite network knowledge bases [63], Re-feeding prior dialogue turns [105], Discounted back-tracking [36], MATRIX running textual memory [98], DeCoAgent persistent long-term storage [37], CompeteAI daybook and comment histories [85]
Widely used techniques Context window memory [68, 130, 62, 153, 86, 129, 132, 61, 118, 81, 70, 55], Conversational and dialogue history [97, 65, 81, 120, 69, 83, 77], Reflexion-style memory systems [67, 102, 13, 111], RAG methods [75, 37, 83, 55], Working memory and scratchpad [67, 49, 124], Long-term episodic memory [68, 123], Short-term memory optimization [111, 81, 73, 78]
@@
The evolution from single-agent to multiagent LLM systems represents a paradigmatic shift toward distributed intelligence architectures that include collaboration, specialization, and coordination to address complex real-world challenges. Multi-agent frameworks represent a significant evolution in AI capabilities, enabling complex problem solving through collaboration, specialization, and coordination. This shift has expanded new frontiers across diverse domains where collective intelligence offers advantages in robustness, scalability, and adaptability [35, 42, 25, 95, 36].
@@
Healthcare. It is among the most impactful domains that benefit from multiagent LLM-based systems. These systems improve diagnostic accuracy through expert role assignment and collaborative reasoning [39, 77, 76] and enhance electronic health record (EHR) processing through coordinated agent interaction [75]. Applications also include triage decision support [38], scenario-based medical training [83], and co-dialogue agents for clinical support [77]. Such systems enable distributed medical reasoning, integrating diverse knowledge modalities while improving interpretability and human-AI alignment.
@@
Software engineering and scientific discovery. Multi-agent systems are also capable of optimizing complex workflows in the engineering domain. Agents collaboratively design hardware systems [74], co-develop AI pipelines [121], and solve engineering problems through shared knowledge and dynamic role allocation [34]. In computational biology and chemistry, agents combine domain-specific tools with LLM reasoning for protein discovery [69], molecular analysis [78], and automated lab planning [56], highlighting the benefits of combining symbolic, statistical and simulation-based reasoning between agents.
@@
Robotics and embodied AI. Multi-agent LLMs coordinate robotic units in both virtual and physical environments. Agents collaboratively plan multirobot tasks [61], generate adaptive driving simulations [120], and train embodied agents using parallel text world simulations [123]. Expanding on these capabilities, LLM-based multiagent systems like CoELA demonstrate how language interfaces can enable decentralized planning, communication, and cooperation in complex embodied environments [60].
@@
Urban planning and telecommunications. Domains such as urban planning and communication are among the few unexplored domains that have also adopted a multi-agent LLM system. Coordinated agents manage large-scale network infrastructures [73] and satellite communication systems using expert mixtures [63], demonstrating scalability in mission critical and real-time environments. Moreover, the multi-agent approach improves specialization and fault tolerance [95], promotes cumulative problem solving [36], and allows systems that better reflect human collaborative dynamics [19, 129]. As these systems advance in capability and complexity, they are set to transform the way intelligent agents engage in various research domains and real-world applications.

### v2

@@ 4.2 Open-sourced LLMs for agentic applications
Domain Single-Agent Multi-Agent (Single + Multi) Key Differences
Interactive and Embodied Environments [45, 99, 56, 118, 46, 13, 104, 25, 36] MineDojo (Minecraft), MarioAI, ALFWorld SMAC, Overcooked-AI, DeCoAgent (Smart Contracts) AI2THOR, ROS 2, Gazebo, Clearpath Husky Open-source stacks; simulation & robotics; embodied/multimodal interaction; native multi-agent tasks; RL & perception–action loops
Code Generation, API Use, and System-Level Integration [34, 105, 33, 37, 66, 74, 63, 131, 50, 71, 110] Code Interpreters, Copilot, Excel, Power BI, Jupyter AI, Chapyter, CoML, RDKit, Scikit-learn, ChatEDA (OpenROAD), Speechly AutoGen RapidAPI Open-source & proprietary; code execution; API orchestration; system/engineering integration
Knowledge Grounding, Web Search, and Structured Retrieval [34, 63, 43, 88, 37, 98, 51, 38, 54] Bing Search API, Google Search, DuckDuckGo, Wikipedia API, PubMed, UMLS, ESI Handbook, FinBERT, MedRAG, ChemCrow (RoboRXN) Open-source & proprietary mix; textual/structured retrieval; domain-specialized; RAG-ready
This domain is further enriched by multimodal and architectural developments. Vision language assistants such as LLaVA-7B and the more recent LLaVA-v1.6-mistral-7B combine a CLIP-style image encoder with a chat-based LLM, allowing agent capabilities in embodied environments and GUI-based interfaces [127, 60]. CogVLM-17B expands this functionality with enhanced visual reasoning and understanding capacity [57]. In the realms of perception and generative synthesis tasks, agents commonly use OpenAI CLIP, Stable Diffusion XL for high-resolution image generation, and VQ-GAN to produce latent image tokenization [56]. These components are often combined with the LLaMA-2 models in interactive applications. These model ecosystems establish the foundational open source landscape on which LLM agent research is currently based.
A critical comparison of proprietary versus open-source models reveals divergent evolutionary trajectories in LLM agent research. Although proprietary models such as GPT-4 and Claude demonstrate superior performance on complex reasoning benchmarks, their widespread deployment involves notable trade-offs. First, the cost of inference remains prohibitively high for large-scale or real-time applications. Second, the closed-source nature of GPT models restricts model transparency and limits fine-tuning flexibility compared to open-source alternatives. Furthermore, reliance on cloud-based APIs raises privacy and security concerns for sensitive domains such as healthcare and finance, where data cannot be externally shared. Recent studies also indicate that domain-specific tasks often require task-adaptive fine-tuning to achieve competitive accuracy. In this regard, open-source models such as Mistral-7B, DeepSeek-V2.5, and LLaMA-2-70B, when fine-tuned effectively, can achieve between 60% and 95% of GPT-4 and other models’ performance while being significantly more cost-efficient [93, 63, 72, 50]. As a result, hybrid model architectures are increasingly adopted, where proprietary models are used as a high-level supervisor for data generation or validation, while smaller open-source models manage routine inference and domain-specific downstream tasks.
@@ 5.1 Usage of tools across knowledge grounding, web search, and structured retrieval
In addition, these tools are often paired with structured external knowledge sources. In particular, agents operate in domain-specific settings. For example, biomedical agents are linked to PubMed and UMLS knowledge bases, one also found using the Emergency Severity Index (ESI) handbook [37] for access to specialized medical information, FinBERT to classify sentiment of financial texts, while the Wikipedia Web API is a widely used source for general encyclopedic information [98, 43, 37, 88, 51].
@@ 5.2 Usage of tools across code generation, API use, and system-level integration
A widely used tool in agentic systems is the diverse types of Code Interpreter developed in a number of studies [34, 105]. This is particularly relevant for data science agents, which are evaluated for their integration with platforms and tools such as Excel, Copilot, Power BI, Jupyter AI, Chapyter, and CoML [105]. Some studies also incorporate AutoGen, a framework designed to build agentic AI systems through multi-agent interactions and improved LLM inference capabilities [33, 37, 66, 74]. These agents often produce code to interact with file systems [63], command-line utilities [63], and domain-specific libraries, such as RDKit for chemical informatics [131] and Scikit-learn for machine learning. Qin et al. found that using ChatGPT’s function calling mechanism, LLMs can interact with 16,464 RESTful APIs sourced from RapidAPI by utilizing a neural API retriever and structured API documentation [50].
In the EDA sector, the ChatEDA agent demonstrates this capability through its Python wrapper-based interaction with the OpenROAD platform [71]. Cuadra et al., in their study, used Speechly to support the speech-to-text task for the purpose of entering health data [110].
@@ 5.3 Usage of tools across interactive and embodied environments
To investigate complex and interactive behaviors, agents are often employed in simulated or real-world environments. They function as tools that offer detailed, state-dependent feedback in response to agent interactions. Zhu et al. [45] employed AI2THOR, a simulation platform that supports realistic 3D environments with physical interactions and dynamic visual states, for the purpose of assessing embodied agents within household environments.
In human-robot interaction, Frering et al. [99] utilized ROS 2 and the Gazebo simulator with a model like the Clearpath Husky. Interactive environments used in agentic research include virtual platforms like game worlds such as Minecraft (via MineDojo) [56, 118], the MarioAI platform [46], and text-based simulations such as ALFWorld [13]. In multi-agent contexts, popular benchmarks include the StarCraft Multi-Agent Challenge (SMAC) and the collaborative game Overcooked-AI [104, 25]. An especially innovative example can be seen in the DeCoAgent framework, where LLM agents autonomously coordinate with the help of smart contracts, overcoming limitations of static, closed multi-agent environments [36].
@@ 6 Frameworks for building LLM agents
This section systematically reviews the frameworks developed for building LLM agents, highlighting their architectural designs, core capabilities, and operational domains. Categorizing existing solutions and analyzing their design principles provides a foundational lens for understanding how LLM agents are constructed and deployed in diverse scenarios.
@@ 6.1 Basic architecture of an LLM agent
Profiling. The profiling module defines an agent’s operational persona (e.g., developer, advisor, or task-specific role), conditioning its behavior and role policy through static profiles defined by experts or dynamic generative mechanisms [21, 138]. Static profiles encode domain knowledge and role-specific behaviors, while dynamic profiles simulate human-like variability by generating diverse agent personas through prompt engineering or parameter sampling [102]. These profiles may include demographic data, personality traits, and social relationships, significantly influencing downstream decisions in memory retrieval, planning strategies, and action selection.
Memory. Memory enables agents to maintain context across interactions through short-term (prompt-based) and long-term (externally stored) forms. Short-term memory supports in-the-moment reasoning by storing dialogue history and environmental signals, but is limited by LLM context windows [30]. Long-term memory captures reusable skills, patterns, or tools from past interactions. Memory formats vary from natural language, embeddings, databases (e.g., SQL), to structured lists, each chosen based on task requirements. Common operations include reading (prioritized by recency, relevance, and importance), writing (handling duplication and overflow), and reflection (summarizing past experiences into high-level insights).
Planning. Planning modules decompose complex tasks using strategies like CoT [4] and Tree-of-Thought (ToT) [139]. Planning can occur with or without feedback. Feedback-free strategies often use stepwise prompting (e.g., CoT, ToT). Feedback-based iteration allows agents to dynamically adapt plans using signals from the environment, human input, or memory reflections.
@@
Together, these components enable LLMs to function as autonomous agents, reasoning, remembering, planning, and acting in open-ended and evolving tasks. This modular architecture is foundational to both single-agent and multi-agent systems discussed in Section 7.
@@ 6.1.1 Single-agent LLM system
LLM is the core of an agent, which allows task planning and decision making based on current inputs, memory, and feedback [141]. In addition, a objective defines the target goal and guides how the agent breaks down complex tasks to formulate a strategy. Memory is a dynamic storehouse of past interactions and relevant contextual information, enabling the agent to adapt its behavior thoughtfully [142, 143]. Action refers to the operational capabilities of the agent. It executes commands, interacts with tools, or communicates outputs [14]; and Rethink facilitates reflective learning by evaluating previous actions and environmental feedback to inform future decisions [15].
@@
Figure 5: A general overview of a multi-agent LLM system. Here, three agents operate within a multimodal environment where they act, generate results, and exchange feedback. Each agent is equipped with internal modules (brain), including memory, reasoning, and planning, that guide its behavior. Through collaborative communication, agents perceive the environment, coordinate strategies, and ultimately take action.
@@ 6.1.2 Multi-agent LLM systems
Multi-agent LLM systems are designed for coordinated collaboration, where multiple agents communicate, adapt, and solve problems together (Figure 5). A description of how multiple agents observe, interact with, and adjust to different environments to work together toward shared objectives is given to highlight the foundation for effective multi-agent collaboration.
@@
Communication structures. In this layer, three major paradigms are prevalent for agent communication, including cooperative (to work together toward shared goals), debate (to argue opposing views to reach consensus), and competitive (to pursue individual objectives). Communication can be centralized, decentralized, or shared memory-based, where agents publish and subscribe to a shared message pool.
@@ 6.2 Common LLM agent frameworks
Several frameworks have been adopted to implement LLM agents, facilitating reasoning, decision making, memory management, and action execution in single-agent and multi-agent contexts. Figure 6 presents a categorized overview of these frameworks based on their application in single-agent or multi-agent systems.
AgentsGenerative Agent Multi: [61] AIDE Single: [86] BDI Single: [99, 124] MetaGPT Multi: [105, 91, 35] CAMEL Multi: [35, 82, 22, 102, 124] AutoGen Multi: [37, 82, 66, 75, 74] LangChain Single: [84, 112, 115, 61, 54] Multi: [102] Reflexion Single: [87, 64, 98, 109, 111, 65, 13, 96, 49] ReAct Single: [13, 96, 87, 23, 64, 65, 108, 98, 52]        [106, 49, 109, 77, 112, 54]
Figure 6: Common LLM agent frameworks used in single- and multi-agent LLMs
Among these, ReAct and Reflexion are the prominently utilized frameworks within single-agent LLM architectures. The selection between ReAct and Reflexion in LLM agents depends on the nature of the task. ReAct, by interleaving structured reasoning with action, such as planning or tool use, facilitates effective and context-aware execution. In contrast, Reflexion enables agents to retrospectively evaluate their prior outputs, learn from feedback, and refine subsequent decisions. Integrating both mechanisms allows for a balance between immediate operational efficiency and iterative self-improvement in complex tasks. AutoGen and CAMEL are widely employed in multi-agent environments. AutoGen is designed for building customizable, complex workflows that emphasize structured role coordination and task orchestration. CAMEL, on the other hand, focuses on autonomous role-playing and goal-driven collaboration, making it especially suited for dialogue-intensive or simulation-based environments. Furthermore, LangChain is also recognized as a versatile framework, applied across single-agent and even in multi-agent paradigms. While it offers a flexible, modular infrastructure for LLM applications, its relatively unopinionated design can limit its applicability for certain tasks, as it places the responsibility for key architectural decisions largely on developers.
@@ 6.3.1 Single-Agent LLM Systems’ Application Domain
Healthcare. LLM agents in healthcare have shown promise in tasks ranging from personalized consultation to domain-specific scientific analysis. Conversational agents provide personalized health advice and support elderly care through natural interfaces [77, 79]. Others optimize public health policies using reinforcement learning (RL) [103] or operate in real-world clinical settings [78]. More specialized agents help with cancer protein analysis and biomedical research tasks [76], underscoring the ability of single agent systems to operate in high-stakes and knowledge-intensive settings.
@@
General-purpose systems. A growing body of work seeks to develop general-purpose agent frameworks that empower LLMs with reasoning, planning, and tool use capabilities [50, 108, 63, 51, 44]. These systems support multitask environments [87], enhance collaboration with humans [98], and enable long-horizon planning through world knowledge and instruction tuning [48, 96]. Several benchmarks have been proposed to assess performance in data science [105], safety [106], and reasoning tasks [47, 141], while architectures address self-improvement [94] and causal reasoning [122]. RL paradigms adapted to LLM further boost agent autonomy [65, 13, 125].
@@ 6.3.2 Multi-Agent LLM systems’ application domain
The evolution from single-agent to multi-agent LLM systems represents a paradigmatic shift toward distributed intelligence architectures that include collaboration, specialization, and coordination to address complex real-world challenges. Multi-agent frameworks represent a significant evolution in AI capabilities, enabling complex problem-solving through collaboration, specialization, and coordination. This shift has expanded new frontiers in diverse domains where collective intelligence offers advantages in robustness, scalability, and adaptability [34, 41, 25, 91, 35].
@@
Healthcare. It is among the most impactful domains that benefit from multi-agent LLM-based systems. These systems improve diagnostic accuracy through expert role assignment and collaborative reasoning [38, 74, 73] and enhance electronic health record (EHR) processing through coordinated agent interaction [72]. Applications also include triage decision support [37], scenario-based medical training [80], and co-dialogue agents for clinical support [74]. Such systems enable distributed medical reasoning, integrating diverse knowledge modalities while improving interpretability and human-AI alignment.
@@
Software engineering and scientific discovery. Multi-agent systems are also capable of optimizing complex workflows in the engineering domain. Agents collaboratively design hardware systems [71], co-develop AI pipelines [117], and solve engineering problems through shared knowledge and dynamic role allocation [33]. In computational biology and chemistry, agents combine domain-specific tools with LLM reasoning for protein discovery [66], molecular analysis [75], and automated lab planning [54], highlighting the benefits of combining symbolic, statistical, and simulation-based reasoning between agents.
@@
Robotics and embodied AI. Multi-agent LLMs coordinate robotic units in both virtual and physical environments. Agents collaboratively plan multi-robot tasks [59], generate adaptive driving simulations [116], and train embodied agents using parallel text world simulations [119]. Expanding on these capabilities, LLM-based multi-agent systems like CoELA demonstrate how language interfaces can enable decentralized planning, communication, and cooperation in complex embodied environments [58].
@@
Urban planning and telecommunications. Domains such as urban planning and communication are among the few unexplored domains that have also adopted a multi-agent LLM system. Coordinated agents manage large-scale network infrastructures [70] and satellite communication systems using expert mixtures [61], demonstrating scalability in mission-critical and real-time environments. Moreover, the multi-agent approach improves specialization and fault tolerance [91], promotes cumulative problem solving [35], and allows systems that better reflect human collaborative dynamics [19, 124]. As these systems advance in capability and complexity, they are set to transform the way intelligent agents engage in various research domains and real-world applications.